OpenStudy (anonymous):

Cookies (c) and bottles of water (w) are sold at a snack bar. The cost of 4 cookies and 2 bottles of water is $3.90 before tax. The cost of 3 cookies and 3 bottles of water is $4.05 before tax.

OpenStudy (anonymous):

here

OpenStudy (anonymous):

so wat r u asking

OpenStudy (anonymous):

What's your question...?

OpenStudy (anonymous):

where is the rest of the problem?

OpenStudy (anonymous):

oh, sorry hold on

OpenStudy (anonymous):

thats fine

OpenStudy (anonymous):

a) Write the systems of equations that represent the information b) Use the system of equations to determine the cost of 1 cookie and the cost of 1 bottle of water before tax

OpenStudy (anonymous):

okay let me see

OpenStudy (anonymous):

@Mathbreaker

OpenStudy (anonymous):

take ur time

OpenStudy (anonymous):

so its 4c + 2w = 3.9 3c + 3w = 4.05 solve these equations to get teh answer :) Simultaneous equations :)

OpenStudy (anonymous):

okay 4c+2w=$3.90

OpenStudy (anonymous):

c4+2w=3.90 w~~0.5 (3.9-c^4)

OpenStudy (anonymous):

c = 0.6 and w = 0.75 just in case youre confused :)

OpenStudy (anonymous):

Therefor w is simply fifty cents

OpenStudy (anonymous):

ya I kinda was lol, I kinda suck at math

OpenStudy (anonymous):

thx guys, so jsut to be sure the answer is $.50

OpenStudy (anonymous):

*just

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=c4%2B2w%3D3.90+ You can check here

OpenStudy (anonymous):

Haha its all about practice i bet you dont :) Look up simultaneous answers if you want help with that No i dont think so, i solved it on wolfram alpha and got http://www.wolframalpha.com/input/?i=4c+%2B+2w+%3D+3.9%3B+3c+%2B+3w+%3D+4.05 Im too lazy to do the math but i think tits 0.6 and 0.75

OpenStudy (anonymous):

ok thx!

OpenStudy (anonymous):

Yeah wolfram's my rock.

OpenStudy (anonymous):

c4+2w=3.90

OpenStudy (anonymous):

Emilyowl, its not c^4.. its 4c + 2w = 3.9 this isnt an exponential function its simply simultaneous :)

OpenStudy (anonymous):

its not c4 guys! Its 4*c + 2*w = 3.9 c4 would be c * c*c*c.... thats weird :P

OpenStudy (anonymous):

Ill give you a final answer 4c + 2w = 3.9 3c + 3w = 4.05 Solving them simultaneous according to this link: http://www.wolframalpha.com/input/?i=4c+%2B+2w+%3D+3.9%3B+3c+%2B+3w+%3D+4.05 Gives me c = 0.6, w = 0.75. And i am absolutely certain its right.

OpenStudy (anonymous):

If you were to use c as a real number hypothetically, then the problem would be exponential.

OpenStudy (anonymous):

What are you implying emilyowl? That makes no sense. Its not exponential. its a system of equations, thats simultaneous :|

OpenStudy (anonymous):

I'm not implying anything, what I said is what I've learned. But it doesn't really matter anymore so okay.

OpenStudy (anonymous):

haha alright im confused about why you put c^4, but lets leave it at this haha